OpenStudy (anonymous):

A man is on a diving board that is 50 feet above the ground. Suppose the man's starting upward velocity is 5 ft/s. Use 0 = -16t^2 + vt + c, where c is the starting height. Find the number of seconds t before he hits the water. Round to the nearest tenth of a second.

OpenStudy (johnweldon1993):

We have each variable we need \[\large −16t^2+vt+c=0\] Plugging in c = 15 and v = 5 we would have \[\large −16t^2+5t+50=0\] From there, we just solve the quadratic

OpenStudy (anonymous):

@johnweldon1993 So I just solve -16t^2 + 5t + 50 = 0?

OpenStudy (johnweldon1993):

mmhmm :)

OpenStudy (anonymous):

Ok thank you!

OpenStudy (johnweldon1993):

Not a problem!